JJ's real estate training school, in Richmond, stocks workbooks with the following characteristics: 

Demand  19,100 units/year

Ordering cost $253/order 

Holding cost $33/unit/year 


The annual holding costs for the workbooks is _______ (round your response to the nearest whole number)
Answer #1

Economic Order Quantity (EOQ) = Square root of (2 x S x D / H)

S = Set up costs (ordering costs)

D = Annual demand

H = Holding costs per unit per year

EOQ = Square root of of (2 x 253 x 19100 / 33)

EOQ = 541 units

Annual Holding costs = 1/2 * EOQ * Holding costs per unit per year = 1/2 x 541 x 33 = $ 8927